OpenStudy (anonymous):

How do I differentiate for dV/dT L(R^2arccos(1-h/r)) -(R-h)(square root of 2hr-r^2)

OpenStudy (anonymous):

\[L(R^2\cos^{-1}(1-h/r) - (R-h)(\sqrt{2hR-h^2)} \] Help differentiating this equation please!

OpenStudy (anonymous):

The answer is \[2L \sqrt{h(2R-h^2)}\] but I don't understand how to get that

OpenStudy (anonymous):

Well, this isn't an equation. I take it this is supposed to be a function, but it's not clear what the independent variable is in this function, making it impossible to know how it is supposed to be differentiated. Can you reproduce the question you are asked, completely?